LBTO:Juan CArlos 9remote), Alessandro (telescope)

Arcetri: Enrico(remote)


Installation the upper clamp(bracket) to keep the OCAM in posiiton.

Rotation test of the pupil wobble and camera lens range.

After the installaiton of the upper clamp the pupils were shifted to the bottom edge of the CCD and it was required to move them up. Alessandro managed to move them up almost ot the center of the CCD releasing the laterla clamps. once the tighten the clamps, the rotaiton check was performed showing that a more stable wobble in both axis. Also, the range of the camera lens are inside. Digital pupil shift were applied and it still requires a fine alignment.


9:18am Alessandro at the telescope and preparing the tools and the new bracket for installaiton.

9:40am LBTI request to have the telescope at 60 degrees. The camera requires to be checked.

10:20 telescope Horizon. Installaiton of the bracket.


10:44 telescope zenith

setFLAt->638 modes.


Mon Nov 15 17:54:20.787 2021 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[3535] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:11,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Mon Nov 15 17:54:20.804 2021 - WFS source set to: FL

cancel preset.

Mon Nov 15 18:03:20.889 2021 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[3537] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:11,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

load pupil: 20211110_215148

First pupil checks.

The pupils are close the botton of the frame.

pupils shift from the puopil 2021110_215148


spilt y=13

loop close at 341.pupil error <0.1, runnig wobble backward

at 300 degres there is a shift in the pupil illumination to the left pupils.

three times and it is the same behaviour.

Skip frames when it reaches a non illumination.

Checked hexapod in postions. Check offload.

telescope to horizon to move the pupils a bit up.

Alessandro goes up to reposition the pupils more at the center. With the top clmap the pupils were moved backward to the end of the CCD frame.


After the adjustment



adjustment of the pupils move up.

TN: 20211115_200642

move to zenith


loaded old pupils: 20210510_221455 #54

spilt x=0, y=1

telescoe at zenith

set flat->638

deploy argos cal unit

Mon Nov 15 20:21:32.761 2021 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[3540] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:11,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)
Mon Nov 15 20:2

Setting pupil: 20210510_221455

start backward


pupil shift.

1:44pM move backward

same behaviour of loosing the PSF.

REstart Adsec adsc-stop/start


setflat->638 modes

Mon Nov 15 21:00:30.973 2021 aos.PresetAO{FLAO}.started - PresetAO {FLAO}[3542] (aomode:ACE-AO,focst:bentGregorianFront,instr:LUCIFER,wfs:FLAOWFS,sox:0[0],soy:0,rox:0[0],roy:0,mag:11,cindex:0,,elev:nan,rotang:nan,grvang:nan,r0:0,skybr:nan,windsp:nan,winddir:nan)

offload not running ..? Activate manually from the adsec control gui

Discussing with Enrico...

Running test with the pupil:
  • 20211115_204005

  • good.
  • good



start to saturate almost at the end.

shift one pixel Y (-1) and X+1

no working

shift x (1) y(1) from 204005 ->20211115_220648





Still requires from this point two pixels down fior the y axis
  • y->-2
  • X is in range.
  • Stable movements in both axis
build a pupil shift from the pupil 220648 with x=0,y=-2

The split requires

one pixel in X->from 4 to 5

to try tomorrow.



-- JuanCarlosGuerra - 15 Nov 2021
Topic revision: r3 - 16 Nov 2021, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback